OpenStudy (anonymous):

PLEASE HELP!!! Identify a function f and a point x = a for which the given limit represents f '(a). lim as x approaches -4 ((sqrt. 29+x) -5) / (x+4)

OpenStudy (anonymous):

there are two parts: f(x)= ? a= ?

zepdrix (zepdrix):

Ummm, here is a form of the limit definition of the derivative that is less frequently used, \[\lim_{x \rightarrow a}\frac{ f(x)-f(a) }{ x-a }\] Now let's compare that to the equation we were given, \[f'(a) =\lim_{x \rightarrow -4}\frac{ \sqrt {29+x} -5}{ x+4 }\] Hmmm they look kind of similar, can you see the f(x) and a? :)

OpenStudy (anonymous):

wow. i totally didn't notice. i was doing it all by hand but i still couldn't come up with it. but thanks!!!

zepdrix (zepdrix):

^^